# TrajFlow: Multi-modal Motion Prediction via Flow Matching

### v1

@@ III-A Preliminaries
with the velociy field $V^{t}$ guiding the probability flow from noise to data distribution. Although $V^{t}$ depend on $t$ in general, linear flow assumes a constant velocity for simplicity. During training, we sample $t$ from a predefined distribution, e.g., uniform, and train a neural network $v_{\theta}$ to approximate $V^{t}$ by minimizing
@@ III-B2 Trajectory-Space Flow Matching Decoder
We adopt the query-based attention decoder design [13, 17, 18, 19, 20, 21, 22, 23, 24], as shown in Fig. 2. To enable multi-modal trajectory prediction, we generate $N_{q}$ query tokens per agent. The decoder then takes as input the flow time $t$ and the noise-corrupted trajectory $Y^{t}$. For the $i$-th query, its initial token is $Q_{i}^{0}=\mathrm{MLP}\left(\texttt{cat}(\mathrm{MLP}(Y^{t}),\mathrm{PT}_{t},\mathrm{PQ}_{i})\right)\in\mathbb{R}^{D}$, where $\mathrm{PT}_{t}$ is the flow time positional embedding, and $\mathrm{PQ}_{i}$ is a learnable positional embedding. The full set of initial queries is given by $Q^{0}=\texttt{cat}(Q_{1}^{0},\dots,Q_{N_{q}}^{0})\in\mathbb{R}^{N_{q}\times D}$. Next, the query tokens go through multiple layers of inter-query self-attention and query-to-context cross-attention, defined at the $j$-th layer as
@@ III-B3 Loss Functions
Before computing the $\arg\min$ operator, we apply Non-Maximum Suppression (NMS) based on trajectory logits, following [23]. We then employ a regression loss to encourage the best trajectory to closely match the ground truth and a classification loss to maximize its score while minimizing the scores of others. Specifically, we define,
@@
Learning-to-Lank Loss. The confidence scores $\hat{S}$ of predicted trajectories are not necessarily calibrated, i.e., a higher score does not always correspond to better accuracy. To promote calibration, we introduce a ranking loss based on the Plackett-Luce (PL) distribution [40, 41]. To this end, we add a prediction head in $F_{\theta}$ to assign ranking scores $\{r_{k}\mid k=1,\dots,N_{q}\}$, where $r_{k}$ represents the preference for $k$-th prediction for the ego agent. We define the ground-truth ranking as $\sigma=\arg\mathrm{sort}([d_{1},\dots,d_{N_{q}}])$, where $d_{k}=\|\hat{Y}^{1}_{k}-Y^{1}\|_{2}^{2}$ is the displacement error. Under the PL model parameterized by $\{r_{k}\}$, the probability of the ground-truth ranking $\sigma$ is given by:
@@
$L=L_{\text{flow}}+L_{\text{cls}}+L_{\text{rank}}.$ (8)
@@ III-B5 Inference
As described in Sec. III-B, each agent generates $N_{q}$ trajectory outputs to capture the multi-modal nature of dynamic motion. Prior works [19, 20, 23, 24] suggest using a larger $N_{q}$ to better cover diverse motion patterns and reflect real-world dynamics. However, $N_{q}$ often exceeds $K$, the required number of trajectories for evaluation. To address this, we apply the commonly used Non-Maximum Suppression (NMS) approach to select $K$ trajectories per agent. NMS leverages confidence scores $\mathrm{sigmoid}(S_{k})$ and an endpoint distance threshold to suppress redundant trajectories, ensuring that the final selection remains both representative and diverse. We use the hyper-parameters from [23] to implement NMS.
Figure 3: Qualitative results. We visualize trajectory predictions and their normalized confidence scores. The ground truth is shown in red with flagged endpoints, the most confident prediction is marked in blue, and other agents are shown in gray.
@@ IV-A2 Metrics
We evaluate performance using multiple metrics, including minimum Average Displacement Error (minADE), minimum Final Displacement Error (minFDE), and Miss Rate. minADE computes the average $\ell_{2}$ distance between the ground-truth trajectory and the closest of $K=6$ predictions across all timesteps, while minFDE measures the final timestep error, emphasizing long-term accuracy. Miss Rate quantifies the proportion of cases where minFDE exceeds a threshold, indicating significant deviation. Additionally, we use mean Average Precision (mAP) to assess multi-modal prediction precision. For further details, please refer to the official leaderboard [8].
@@ IV-C Quantitative Results
We present the main results in Tab. I on the marginal prediction testing set of WOMD. Our method achieves competitive performance, outperforming all baselines in terms of minADE, minFDE, and Miss Rate. It attains the second-highest mAP and soft mAP, trailing only behind BeTop by no more than $0.005$, while still surpassing other baselines by a non-trivial margin. Notably, these results were obtained using a single model for forecasting, without employing any ensemble techniques. To ensure a fair comparison, we exclude methods that utilize additional data, such as LIDAR or camera inputs. Additionally, we compare our approach with a model variant that does not incorporate the PL ranking loss, as defined in Eq. 7. The results clearly demonstrate that the learning-to-rank regularization loss effectively enhances the model’s capacity across a wide range of metrics.
@@ IV-E Ablation Studies
To verify the introduced self-conditioning mechanism and the PL ranking loss, we conduct ablation studies by training various model variants using 20% of the training set and evaluating them on the validation set. As shown in Tab. II, self-conditioning significantly improves performance as the number of ODE sampling steps increases (see Sec. III-A for details on ODE sampling). In contrast, without self-conditioning, the model’s generalization degrades with more sampling steps due to the overfitting issue discussed in Sec. III-B4. Additionally, the PL ranking loss consistently enhances performance, particularly in mAP, showcasing its positive impact on learning calibrated confidence scores.
@@ A-A1 Training and Inference Pipeline
11:         $\theta\leftarrow\theta-\eta\nabla_{\theta}L(\check{S},\check{Y}^{1},\check{r},Y^{1})$triangleright Loss in Eq. 8
12:     end if
13:     $\hat{Y}^{1},S,r\leftarrow F_{\theta}(Y^{t},C,t)$
14:     $\theta\leftarrow\theta-\eta\nabla_{\theta}L(S,\hat{Y}^{1},r,Y^{1})$triangleright Loss in Eq. 8
15:until converged
@@ A-A2 Model Configurations
$L_{\text{rank}}$ 1.0
As shown in Table III, TrajFlow primarily comprises a context encoder, flow matching decoder, and prediction heads. The context encoder has 6 layers, an embedding dimension of 256, and 8 attention heads. The flow matching decoder also has 6 layers but with a larger embedding dimension of 512 and 8 attention heads. The prediction heads for trajectory, classification, and ranking each use a 3-layer MLP. Training is conducted using the AdamW optimizer [44] with $\beta_{1}=0.9$, $\beta_{2}=0.999$, a linear decay learning rate scheduler over 22/24/26/28 epochs, 40 total epochs, a batch size of 80, a peak learning rate of $1\times 10^{-4}$, and a weight decay of 0.01. Loss weights for flow matching denoising ($L_{\text{flow}}$), classification ($L_{\text{cls}}$), and ranking ($L_{\text{rank}}$) are all set to 1.0. To adapt to the denoising objective, we map trajectory data (X-Y coordinates) to the range $[-1,1]$ using square-root-based normalization. For each coordinate, we subtract an offset and multiply by a scaling coefficient to ensure consistent scaling across the dataset based on predefined metadata parameters. We collect statistics from the training set to ensure that over 99.9% of data points can be reversibly mapped to the target range without loss. This normalization aligns with previous work on diffusion and flow-matching generative models [47, 48, 32, 49, 30, 50], which aims improve training stability.
A-B Additional Experimental Results
@@
TrajFlow (Ours) 0.9403 2.1685 0.4177 0.2434 0.2491
Figure 4: Additional qualitative results. We visualize trajectory predictions with normalized confidence scores: ground truth in red with flagged endpoints, top prediction in blue, and other agents in gray.
@@ A-B1 More Quantitative Results
Our method outperforms all baselines on the ADE, FDE, and Miss Rate metrics on the standard split of the testing dataset. It ranks second only to BeTop, with a slight difference in mAP and Soft mAP. On the standard validation set, our method performs strongly in terms of mAP and Soft mAP, indicating a robust capacity for learning uncertainty, while maintaining a very small gap compared to the state-of-the-art in ADE, FDE, and Miss Rate. On the interactive dataset, our method surpasses the baselines in both mAP and Soft mAP. Please note that, due to the lack of open-source resources, we were unable to fully reproduce the results of previous models on the interactive sets; therefore, we compare only against the numbers reported in their papers.
A-B2 More Qualitative Results
Additional qualitative results are presented in Fig. 4 and Fig. 5, providing further examples of both successful and failed cases on the validation datasets. In Fig. 4, we showcase a diverse set of agents—including vehicles, pedestrians, and cyclists—making reasonable predictions. In these cases, at least one of the predicted trajectories closely aligns with the ground truth, while the remaining predictions exhibit diverse and plausible motion patterns. This highlights the model’s ability to capture multi-modality and generate varied hypotheses about future movements. Despite the generally strong performance, the proposed TrajFlow model encounters difficulties in certain challenging scenarios, as illustrated in Fig. 5. These failure cases often involve complex interactions, rare or unseen motion patterns, or ambiguous intent, all of which can mislead the model. We attribute these limitations to a combination of data scarcity in such edge cases and the inherent uncertainty of real-world driving situations, which require a deep understanding of contextual cues and agent intent.

### v2

@@ III-A Preliminaries
with the velocity field $V^{t}$ guiding the probability flow from noise to data distribution. Although $V^{t}$ depends on $t$ in general, linear flow assumes a constant velocity for simplicity. During training, we sample $t$ from a predefined distribution, e.g., uniform, and train a neural network $v_{\theta}$ to approximate $V^{t}$ by minimizing
@@ III-B2 Trajectory-Space Flow Matching Decoder
We adopt the query-based attention decoder design [13, 17, 18, 19, 20, 21, 22, 23, 24], as shown in Fig. 2. To enable multi-modal trajectory prediction, we generate $N_{q}$ query tokens per agent. The decoder then takes as input the flow time $t$ and the noise-corrupted trajectory $Y^{t}$. For the $i$-th query, its initial token is $Q_{i}^{0}=\mathrm{MLP}\left(\texttt{cat}(\mathrm{MLP}(Y^{t}),\mathrm{PT}_{t},\mathrm{PQ}_{i})\right)\in\mathbb{R}^{D}$, where $\mathrm{PT}_{t}$ is the flow time positional embedding, and $\mathrm{PQ}_{i}$ is a learnable positional embedding that projects the context token of the center agent, which is one element of the agent tokens $Z_{a}$. The full set of initial queries is given by $Q^{0}=\texttt{cat}(Q_{1}^{0},\dots,Q_{N_{q}}^{0})\in\mathbb{R}^{N_{q}\times D}$. Next, the query tokens go through multiple layers of inter-query self-attention and query-to-context cross-attention, defined at the $j$-th layer as
@@ III-B3 Loss Functions
Before computing the $\arg\min$ operator, we apply Non-Maximum Suppression (NMS) based on trajectory logits, following [23]. NMS leverages trajectory logits and an endpoint distance threshold to suppress redundant trajectories, generating a fixed number of trajectories from the $N_{q}$ proposals. Its goal is to select trajectories with representative and diverse motion patterns. We then employ a regression loss to encourage the best trajectory to closely match the ground truth and a classification loss to maximize its score while minimizing the scores of others. Specifically, we define,
@@
Learning-to-Rank Loss. The confidence scores $\hat{S}$ of predicted trajectories are not necessarily calibrated, i.e., a higher score does not always correspond to better accuracy. To promote calibration, we introduce a ranking loss based on the Plackett-Luce (PL) distribution [40, 41, 42]. To this end, we add a prediction head in $F_{\theta}$ to assign ranking scores $\{r_{k}\mid k=1,\dots,N_{q}\}$, where $r_{k}$ represents the preference for $k$-th prediction for the ego agent. We define the ground-truth ranking as $\sigma=\arg\mathrm{sort}([d_{1},\dots,d_{N_{q}}])$, where $d_{k}=\|\hat{Y}^{1}_{k}-Y^{1}\|_{2}^{2}$ is the displacement error. Under the PL model parameterized by $\{r_{k}\}$, the probability of the ground-truth ranking $\sigma$ is given by:
@@
$L=L_{\text{flow}}+L_{\text{cls}}+\lambda L_{\text{rank}}.$ (8)
We set $\lambda=0.1$ in our implementation to boost performance.
@@ III-B5 Inference
As described in Sec. III-B, each agent generates $N_{q}$ trajectory outputs to capture the multi-modal nature of dynamic motion. Prior works [19, 20, 23, 24] suggest using a larger $N_{q}$ to better cover diverse motion patterns and reflect real-world dynamics. However, $N_{q}$ often exceeds $K$, the required number of trajectories for evaluation. To address this, we apply the commonly used Non-Maximum Suppression (NMS) approach to select $K$ trajectories per agent. NMS leverages confidence scores $\mathrm{sigmoid}(S_{k})$ and an endpoint distance threshold to suppress redundant trajectories, ensuring that the final selection remains both representative and diverse. We use the hyper-parameters from [23] to implement NMS, following the same pipeline as during training.
Figure 3: Qualitative results. We visualize trajectory predictions (left) and their normalized confidence scores (right) in each plot. Six trajectory predictions are marked with different colors and numbered at their endpoints; the most confident prediction is shown in blue. The ground truth is displayed in red, with a flag indicating its endpoint. Surrounding road elements, such as lanes and sidewalks, are shown as lines and dots. Other surrounding traffic objects are represented as gray boxes.
@@ IV-A2 Metrics
We evaluate performance using multiple metrics, including minimum Average Displacement Error (minADE), minimum Final Displacement Error (minFDE), and Miss Rate. minADE computes the average $\ell_{2}$ distance between the ground-truth trajectory and the closest of $K=6$ predictions across all timesteps, while minFDE measures the final timestep error, emphasizing long-term accuracy. Miss Rate quantifies the proportion of cases where minFDE exceeds a threshold, indicating significant deviation. Additionally, we use mean Average Precision (mAP) to assess multi-modal prediction precision. To compute mAP, ground truth trajectories are grouped by motion type into buckets. For each bucket, predictions are ranked by confidence and labeled true or false positives, counting only the highest-confidence true positive per object while treating others as false positives. Precision-recall curves are computed using interpolated precision, averaging across buckets for the final mAP. Soft mAP is similar but does not penalize multiple matching predictions beyond the highest-confidence one. Both mAP metrics require prediction scores to preserve relative ranking for accurate evaluation. For further details, please refer to the official leaderboard [8].
@@ IV-C Quantitative Results
We present the main results in Tab. I on the marginal prediction testing set of WOMD. Our method achieves competitive performance, outperforming all baselines across all metrics, including minADE, minFDE, Miss Rate, and two mAP metrics. It attains the highest mAP and soft mAP, and surpasses the other baselines by a non-trivial margin. The mAP and soft mAP metrics are the most crucial indicators of model performance according to the WOMD official evaluation protocol [8]. Notably, these results were obtained using a single model for forecasting, without employing any ensemble techniques. To ensure a fair comparison, we exclude methods that utilize additional data, such as LIDAR or camera inputs. Please refer to Tab. IV for results on the joint motion prediction dataset.
@@ IV-E Ablation Studies
To verify the introduced self-conditioning mechanism and the PL ranking loss, we conduct ablation studies by training various model variants using 20% of the training set and evaluating them on the validation set. As shown in Tab. II, self-conditioning significantly improves performance as the number of ODE sampling steps increases (see Sec. III-A for details on ODE sampling). In contrast, without self-conditioning, the model’s generalization degrades with more sampling steps due to the overfitting issue discussed in Sec. III-B4. Additionally, we compare our approach with a model variant that does not incorporate the PL ranking loss, as defined in Eq. 7. The results clearly demonstrate that the learning-to-rank regularization loss effectively enhances the model’s performance across a wide range of metrics, particularly in mAP. Because mAP effectively measures the alignment between trajectory scores and prediction accuracy, its improvement with the PL ranking loss indicates a meaningful enhancement in uncertainty estimation.
@@ A-A1 Training and Inference Pipeline
11:         $l_{s}\leftarrow L(\check{S},\check{Y}^{1},\check{r},Y^{1})$triangleright Loss in Eq. 8
12:     else
13:         $l_{s}\leftarrow 0$
14:     end if
15:     $\hat{Y}^{1},S,r\leftarrow F_{\theta}(Y^{t},C,t)$
16:     $l\leftarrow L(S,\hat{Y}^{1},r,Y^{1})$triangleright Loss in Eq. 8
17:     $\theta\leftarrow\theta-\eta\nabla_{\theta}(l+l_{s})$
18:until converged
@@ A-A2 Model Configurations
$L_{\text{rank}}$ 0.1
As shown in Table III, TrajFlow primarily comprises a context encoder, flow matching decoder, and prediction heads. The context encoder has 6 layers, an embedding dimension of 256, and 8 attention heads. The flow matching decoder also has 6 layers but with a larger embedding dimension of 512 and 8 attention heads. The prediction heads for trajectory, classification, and ranking each use a 3-layer MLP. Training is conducted using the AdamW optimizer [45] with $\beta_{1}=0.9$, $\beta_{2}=0.999$, a linear decay learning rate scheduler over 22/24/26/28 epochs, 40 total epochs, a batch size of 80, a peak learning rate of $1\times 10^{-4}$, and a weight decay of 0.01. Loss weights for flow matching denoising ($L_{\text{flow}}$) and classification ($L_{\text{cls}}$) are set to 1.0 and the weight for ranking ($L_{\text{rank}}$) is set to 0.1. To adapt to the denoising objective, we map trajectory data (X-Y coordinates) to the range $[-1,1]$ using linear-based normalization. For each coordinate, we subtract an offset and multiply by a scaling coefficient to ensure consistent scaling across the dataset based on predefined metadata parameters. We collect statistics from the training set to ensure that over 99.9% of data points can be reversibly mapped to the target range without loss. This normalization aligns with previous work on diffusion and flow-matching generative models [48, 49, 50, 51, 52, 30, 53, 54], which aims to improve training stability.
@@
TrajFlow (Ours) 0.9273 2.1249 0.4110 0.2464 0.2531
A-B Additional Experimental Results
@@ A-B1 More Quantitative Results
Our method outperforms all baselines across all metrics, including ADE, FDE, Miss Rate, and mAP, on the standard test split of the dataset. On the standard validation set, our method performs strongly in terms of mAP and Soft mAP, indicating a robust capacity for modeling uncertainty, while maintaining a very small gap compared to the state-of-the-art in ADE, FDE, and Miss Rate. On the interactive dataset, our method surpasses the baselines in both mAP and Soft mAP. Please note that, due to the lack of open-source resources, we were unable to fully reproduce the results of previous models on the interactive sets; therefore, we compare only against the numbers reported in their respective papers.
A-B2 Runtime Comparison
TABLE V: Inference runtime comparison.
Method Runtime #Params
Per Agent (ms)$\downarrow$ Total Val. Set (s)$\downarrow$
MTR [19] 19.96 880 65.8M
Ours 16.44 725 53.3M
We present the inference runtime results in Tab. V, comparing MTR [19] and our method on the WOMD validation set, which consists of 44,097 agents, using the same hardware (NVIDIA A100 80GB). Thanks to the multi-shot prediction formulation and few-step sampling, our method achieves relatively fast runtime. Also, our model is slightly smaller than MTR, which further contributes to its faster inference speed.
Figure 4: Additional qualitative results. We visualize trajectory predictions (left) and their normalized confidence scores (right) in each plot. Six trajectory predictions are marked with different colors and numbered at their endpoints; the most confident prediction is shown in blue. The ground truth is displayed in red, with a flag indicating its endpoint. Surrounding road elements, such as lanes and sidewalks, are shown as lines and dots. Other surrounding traffic objects are represented as gray boxes.
A-B3 More Qualitative Results
Additional qualitative results are presented in Fig. 4 and Fig. 5, providing further examples of both successful and failed cases on the validation datasets. In Fig. 4, we showcase a diverse set of agents—including vehicles, pedestrians, and cyclists—making reasonable predictions. In these cases, at least one of the predicted trajectories closely aligns with the ground truth, while the remaining predictions exhibit diverse and plausible motion patterns. This highlights the model’s ability to capture multi-modality and generate varied hypotheses about future movements.
Inaccurate predictions can be attributed to two main factors: (1) the predicted trajectories may deviate significantly from the ground truth in space, leading to prediction errors; and (2) the most accurate trajectories may not receive the highest confidence scores, resulting in calibration errors. The last row of Fig. 4 highlights two special cases where trajectories close to the ground truth are present, but are not ranked highest, indicating a calibration issue. Despite the overall strong performance of the proposed TrajFlow model, it struggles in certain challenging scenarios, as shown in Fig. 5, where predictions spatially diverge from the ground truth. These failure cases often involve complex interactions, rare or unseen motion patterns, or ambiguous intent, all of which can challenge and mislead the model. We attribute such limitations to a combination of data scarcity in these edge-case scenarios and the inherent uncertainty of real-world driving, which demands a nuanced understanding of contextual cues and agent intent.
A-B4 Societal Impact
Our method supports motion prediction for agents in autonomous driving scenarios, contributing to safer navigation and improved decision-making. However, its reliability in edge cases—such as rare behaviors or complex interactions—remains an open question. For safety-critical deployment, these limitations must be addressed through rigorous testing, robust uncertainty estimation, and careful consideration of potential failure modes. Broader societal implications, including fairness across agent types and accountability in decision-making, should also be carefully considered.